### v1

# Two or three things particle physicists (mis)understand about (pre)heating

Abstract
The transition from the end of inflation to a hot, thermal Universe, commonly referred to as (re)heating, is a critical yet often misunderstood phase in early Universe cosmology. This short review aims to provide a comprehensive, conceptually clear, and accessible introduction to the physics of (re)heating, tailored to the particle physics community. We critically examine the standard Boltzmann approach, emphasizing its limitations in capturing the intrinsically non-perturbative and non-linear dynamics that dominate the early stages of energy transfer. These include explosive particle production, inflaton fragmentation, turbulence, and thermalization; phenomena often overlooked in perturbative treatments. We survey a wide range of theoretical tools, from Boltzmann equations to lattice simulations, clarifying when each is applicable and highlighting scenarios where analytic control is still feasible. Special attention is given to model-dependent features such as (pre)heating, the role of fermions, gravitational couplings, and the impact of multifield dynamics. We also discuss exceptional cases, including Starobinsky-like models and instant (pre)heating, where (re)heating proceeds through analytically tractable channels without requiring full non-linear simulations. Ultimately, this review serves both as a practical guide and a cautionary tale, advocating for a more nuanced and physically accurate understanding of this pivotal epoch within the particle physics community.
1 Introduction
Over the past decades, significant progress has been made in understanding the early Universe, with inflation emerging as a cornerstone of modern cosmology. Precision measurements of the cosmic microwave background and the successful predictions of Big Bang nucleosynthesis (BBN) offer compelling evidence for a thermal radiation-dominated phase in the early Universe. Yet, the transition from a cold, inflationary state to the hot Big Bang remains one of the least understood epochs in cosmology. This crucial phase—known as (re)heating [^1] Throughout this review, we adopt the terms (re)heating and (pre)heating, rather than the more commonly used reheating and preheating, to emphasize that, within the standard inflationary scenario, the Universe need not have been in thermal equilibrium before the onset of inflation.—marks the transfer of energy from the inflaton field to the Standard Model (SM) degrees of freedom, enabling the formation of matter and radiation as we know them.
Despite its pivotal role, (re)heating is often oversimplified in the particle physics literature. The commonly used Boltzmann equation (BEQ) framework, which models energy transfer via perturbative inflaton decays or scatterings, fails to account for key features of the process. These include explosive, non-perturbative effects such as parametric and tachyonic resonance, inflaton fragmentation, Bose enhancement, and turbulent thermalization—all of which can dramatically alter the dynamics. In particular, it has been shown that the Boltzmann and Bogoliubov approaches are inequivalent, especially at low-momentum modes where nonadiabatic particle production dominates (see, e.g., Ref. [1]). In reality, the (re)heating phase is a rich multistage process. It begins with a non-perturbative particle production or (pre)heating stage, often requiring numerical lattice simulations to resolve the nonlinear evolution of coupled fields. As the system evolves, it gradually transitions to a more familiar perturbative regime where standard thermalization and decay processes can be described analytically. Both stages are sensitive to the specific inflationary potential, the type of interactions with matter, and the structure of the underlying theory.
The aim of this short review is to provide a pedagogical overview of (re)heating tailored to the particle physics community. Rather than attempting to cover the entire body of literature on the subject (a number of excellent reviews on this subject can already be found in the literature; see, e.g., Refs. [2, 3, 4, 5, 6]), we focus on guiding the reader through the most essential concepts and mechanisms in a schematic and accessible manner. We demystify the limitations of standard perturbative approaches, highlight commonly overlooked non-linear phenomena, and clarify when simplified methods can still yield reliable insights. To this end, we survey a range of theoretical tools—including BEQs, semi-analytical estimates, and lattice simulations—emphasizing the specific conditions under which each framework is applicable. We also explore scenarios in which (re)heating proceeds solely via gravitational or scale-induced couplings, without requiring additional interactions with beyond-the-SM (BSM) fields.
The structure of the manuscript is as follows. In Section 2, we revisit the conventional Boltzmann-based approach to (re)heating and the assumptions behind it. Section 3 explores the limitations of perturbation theory and introduces non-perturbative phenomena such as (pre)heating and inflaton fragmentation. We also highlight how the presence of fermions and multifield dynamics modifies (pre)heating and thermalization. Section 4 then focuses on the exceptional cases where full nonlinear simulations can be avoided and analytic control is still possible. We conclude in Section 5 by summarizing key takeaways and outlining open questions.
2 The poor particle-physics approach
The usual Boltzmann framework for studying (re)heating focuses on tracking the phase-space distribution of particles produced by perturbative decays or scatterings of the inflaton condensate by solving sets of BEQs [7]. This approach enables a systematic analysis of the time evolution of the energy densities associated with the inflaton, radiation, and any other exotic states generated during (re)heating. The Boltzmann framework has been extensively discussed in the literature, particularly within the particle physics community; see, e.g., Refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Here, we provide a concise overview, emphasizing the key assumptions underlying this intrinsically perturbative treatment.
Immediately after inflation, the inflaton field $\phi$ is typically spatially homogeneous, with any residual inhomogeneities or self-interactions considered negligible. The post-inflationary evolution of the Universe is then primarily dictated by the inflaton potential, with minimal contributions from other energy components, until radiation domination takes over. Among the various inflationary models consistent with observations [37], ranging from the seminal Starobinsky model [38, 39, 40, 41] to Higgs inflation [42] and its variants [43, 44, 45, 46, 47] or other universal attractors [48, 49], we focus here for concreteness on a $T$-model of inflation with potential [48]
$V(\phi)=\lambda\,\Lambda^{4}\,\tanh^{n}\left(\frac{|\phi|}{\Lambda}\right)\simeq\lambda\,\Lambda^{4}\times\begin{dcases}\left(\frac{|\phi|}{\Lambda}\right)^{n}&\text{ for }|\phi|\ll\Lambda\,,\\ 1&\text{ for }|\phi|\gg\Lambda\,,\end{dcases}$ (2.1)
smoothly interpolating between a power-law behavior for small field values and a nearly constant plateau at large ones, with $\lambda$ a dimensionless coupling constant and $\Lambda$ an energy scale to be constrained from observations. With the usual chaotic slow-roll conditions [50], this asymptotic plateau allows for an accelerated expansion of the Universe that comes to an end when $\ddot{a}=0$, with $a$ the cosmological scale factor. The subsequent oscillatory evolution of the spatially homogeneous inflaton field around the minimum of its potential (2.1) is customarily described by an effective equation of motion
$\ddot{\phi}+(3\,H+\Gamma_{\phi})\,\dot{\phi}+V_{,\phi}(\phi)=0\,,$ (2.2)
with the dots indicating derivatives with respect to coordinate time $t$, and the commas derivatives with respect to the field $\phi$. Here, $H$ represents the Hubble expansion rate, while $\Gamma_{\phi}$ corresponds to the total inflaton decay rate, which quantifies the energy transfer rate from individual inflaton quanta to the conventional matter and radiation components. This decay rate enters the equation of motion as an additional friction-like term, effectively governing the dissipation of the inflaton energy density over time. In most particle physics approaches, this decay rate is assumed to be either constant or only weakly dependent on time, allowing for a gradual and controlled energy transfer that smoothly connects inflationary (re)heating with the subsequent thermal history of the Universe. The exact form of the decay rate depends on the particular inflaton decay mode, which, in turn, is influenced by the underlying particle-physics model.
In the context of (re)heating, the interactions of the inflaton with both SM and BSM fields can be broadly categorized into two primary mechanisms: direct decay and scatterings. On the one hand, in the decay channel, the inflaton field $\phi$ can transition into pairs of lighter particles via trilinear interactions, such as $\mu\,\phi\,|\varphi|^{2}$ for scalar fields $\varphi$ (e.g., the Higgs boson doublet) or Yukawa-like couplings $y_{\psi}\,\phi\,\overline{\Psi}\,\Psi$ for fermionic fields $\Psi$. These interactions lead to two-body decays, with the coupling parameters $\mu$ and $y_{\psi}$ controlling their respective strengths. On the other hand, scattering processes, which provide an alternative pathway for energy transfer, occur via quartic couplings of the form $g\,\phi^{2}\,|\varphi|^{2}$, allowing 2-to-2 interactions of strength $g$ that populate the thermal bath with SM and BSM particles. Beyond this model-dependent channel, there exists an irreducible gravity-mediated production following from the unavoidable interaction term $\sqrt{-g}\,\mathcal{L}_{\rm grav}\supset-2/M_{P}\,h_{\mu\nu}\,T^{\mu\nu}$ between the graviton field $h_{\mu\nu}$ and the energy-momentum tensor $T_{\mu\nu}$ of all matter fields in the theory [51], with $M_{P}=(8\pi\,G_{N})^{-1/2}\simeq 2.4\times 10^{18}$ GeV the reduced Planck mass. In this case, (re)heating can occur through 2-to-2 scatterings of the inflaton condensate into SM final states [52, 53, 54, 14, 55, 21, 25, 27, 56, 57], provided of course that the available energy density is large enough to partially overcome the involved Planck suppression. Alternatively, gravitational (re)heating can also occur through gravity-mediated decays as in the case of linear non-minimal interactions of the inflaton [58].
Given the above set of interactions and assuming that the bare masses of the decay products are significantly smaller than that of the inflaton condensate, the inflaton’s total decay width, computed using standard Feynman rules techniques, can be approximated as [59, 60]
$\Gamma_{\phi}=\frac{\gamma^{2}_{\text{eff}}}{8\pi}\times\begin{dcases}m_{\phi}(a)&\quad\text{for fermionic decay},\\ m^{-1}_{\phi}(a)&\quad\text{for bosonic decay},\\ \rho_{\phi}/m_{\phi}^{3}&\quad\text{for bosonic scattering},\end{dcases}$ (2.3)
for explicit couplings to matter, and as [24, 27, 56]
$\Gamma_{\phi}\simeq\alpha_{n}\,M_{P}^{5}\left(\frac{\rho_{\phi}}{M_{P}^{4}}\right)^{\frac{5n-2}{2n}},$ (2.4)
for minimal gravitational couplings, with $\alpha_{n}$ encoding the effect of oscillating inflaton condensate. Here, $\rho_{\phi}$ is the inflaton energy density,
$m_{\phi}^{2}(a)\equiv\frac{d^{2}V}{d\phi^{2}}\simeq n\,(n-1)\,\lambda^{\frac{2}{n}}\,\Lambda^{\frac{2\,(4-n)}{n}}\rho_{\phi}(a)^{\frac{n-2}{n}}\simeq m_{\phi}(a_{I})\left(\frac{a_{I}}{a}\right)^{\frac{3(n-2)}{n+2}}$ (2.5)
stands for the inflaton mass in the vicinity of the origin ($|\phi|\ll\Lambda$), and
${\gamma^{2}_{\text{eff}}}=\gamma^{2}\,(n-1)\,(n+2)\left(\frac{\omega}{m_{\phi}}\right)^{q}\,\sum_{j=1}^{\infty}j^{q}\,|\mathcal{P}j|^{2}\,\left\langle\left[1-\left(\frac{2\,m_{\phi}}{j\,\omega}\right)^{2}\mathcal{P}\right]^{q/2}\right\rangle$ (2.6)
[…]
$\phi(t)=\phi_{0}(t)\,\mathcal{P}(t)=\phi_{0}(t)\sum_{j=1}^{\infty}\mathcal{P}_{j}(t)\,e^{-i\,j\,\omega\,t}\,,\hskip 28.45274pt\omega=m_{\phi}\,\sqrt{\frac{n\,\pi}{2n-1}}\,\frac{\Gamma\left(\frac{n+1}{2n}\right)}{\Gamma\left(\frac{1}{2n}\right)}\,,$ (2.7)
where $\mathcal{P}(t)$ is a quasi-periodic, fast-oscillating function, and $\phi_{0}(t)$ is a slowly-varying envelope. Note that for $n=2$, both the inflaton mass and the frequency $\omega$ in this mode decomposition are exactly constant, while displaying an explicit field dependence for $n>2$ which makes them decrease with time.
Under the assumption that the backreaction of the aforementioned relativistic decay products is small, the inflaton phase-space distribution function is given by some sharply peaked or delta function momentum distribution $f_{\phi}(k,\,t)=(2\pi)^{3}\,n_{\phi}(t)\,\delta^{(3)}(\vec{k})$, with $n_{\phi}$ the inflaton number density. The corresponding integrated BEQ for $n_{\phi}$ can be written as [7]
[…]
where $p$ collectively denotes all the final particles produced from inflaton decay or scattering, $d\Pi$’s are the Lorentz-invariant phase space, |¯$|\overline{\mathcal{M}}|^{2}$ is the spin-averaged squared amplitude for the underlying process concerned (with any symmetry factors included), and $\left(1\pm f_{p}\right)$ takes care of Pauli-blocking/stimulated emission effects for the product particles. Clearly, all information from the underlying particle physics model is imprinted in the matrix element $\mathcal{M}$, which is computed perturbatively for a given decay/scattering diagram involving inflaton and SM fields. Now, assuming $i)$ no backreaction producing inflaton condensate from the bath, i.e., |¯$|\overline{\mathcal{M}}|^{2}_{p\to\phi}=0$, and $ii)$ that the Bose enhancement and Pauli blocking effects due to the decay products are negligible (that is, $1\pm f_{p}\simeq 1$), one can arrive at the evolution equation for the inflaton energy density $\rho_{\phi}$, which reads [65]
$\frac{d\rho_{\phi}}{dt}+\frac{6\,n}{2+n}\,H\,\rho_{\phi}=-\frac{2\,n}{2+n}\,\Gamma_{\phi}\,\rho_{\phi}\,.$ (2.9)
For small coupling constants compatible with the radiative instability of the inflaton potential during inflation, the term $H\,\rho_{\phi}$ associated with the cosmological expansion during (re)heating typically dominates over the interaction term $\Gamma_{\phi}\,\rho_{\phi}$. In this regime, the evolution of the inflaton energy density and the associated Hubble rate $H\propto\rho_{\phi}^{1/2}$ evolve as
$\rho_{\phi}(a)\simeq\rho_{\phi}(a_{\text{rh}})\left(\frac{a_{\text{rh}}}{a}\right)^{\frac{6\,n}{n+2}},\quad\quad H(a)\simeq H(a_{\text{rh}})\left(\frac{a_{\text{rh}}}{a}\right)^{\frac{3\,n}{n+2}},$ (2.10)
with $a_{I}$ and $a_{\text{rh}}$ the scale factors at the end of inflation and at the conclusion of the (re)heating stage, respectively. This scaling is associated to an effective equation-of-state [65, 66]
$\langle w\rangle=\frac{\langle p_{\phi}\rangle}{\langle\rho_{\phi}\rangle}=\frac{n-2}{n+2}\,,$ (2.11)
where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes an average over several inflaton oscillations. On the other hand, the evolution of the radiation energy density $\rho_{R}$ is governed by a BEQ of the form [65]
$\frac{d\rho_{R}}{dt}+4\,H\,\rho_{R}=+\frac{2\,n}{2+n}\,\Gamma_{\phi}\,\rho_{\phi}\,,$ (2.12)
which, using Eq. (2.10), admits a solution
[…]
Assuming that the decay products thermalize immediately after production, [^2] For detailed analysis of (perturbative) thermalization see, e.g. Refs. [67, 68, 69, 70]. one can define an instantaneous radiation temperature as
$T(a)=\left(\frac{30\,\rho_{R}(a)}{\pi^{2}\,g_{\star}(a)}\right)^{1/4},$ (2.14)
with $g_{\star}(a)$ the evolving number of relativistic degrees of freedom contributing to the SM energy density. Note that, in general, this temperature can significantly exceed the so-called (re)heating temperature $T_{\text{rh}}\equiv T(a_{\text{rh}})$ marking the onset of radiation domination, $\rho_{R}(a_{\text{rh}})=\rho_{\phi}(a_{\text{rh}})$, facilitating with it key processes such as baryogenesis or the thermalization of weakly-interacting dark matter candidates with the radiation bath [71]. During (re)heating, the SM temperature typically follows as simple scaling
$T(a)\simeq T_{\text{rh}}\left(\frac{a_{\text{rh}}}{a}\right)^{\alpha},$ (2.15)
with $\alpha$ depending on the inflaton potential and the details fo the energy transfer (spin of the daughter particles and annihilation/decay channel). In any case, to ensure the success of BBN, the (re)heating temperature must satisfy $T_{\text{rh}}>T_{\text{BBN}}\simeq 4$ MeV [72, 73, 74, 75, 76, 77].
Figure 1: Summary of different (re)heating scenarios. The black dot corresponds to the case $n=2$, where the inflaton scales as non-relativistic matter and decays into SM particles with a constant decay width, while the black crosses correspond to the alternative scenarios described in the text. The red area in the upper left corner does not give rise to viable (re)heating.
[…]
Figure 2: Evolution of the energy densities as a function of the scale factor $a$ for the inflaton ($\rho_{\phi}$, blue line) and the SM bath $\rho_{R}$, assuming an inflaton decays into fermions (dashed black) or scalars (solid black), for a quartic potential $n=4$.
Figure 2 illustrates the evolution of energy densities as functions of the scale factor $a$ for the inflaton ($\rho_{\phi}$, blue line) and the SM radiation bath ($\rho_{R}$), assuming that the inflaton decays into fermions (dashed black) or scalars (solid black). The results correspond to a quartic inflaton potential, for which $n=4$, and are obtained from a full numerical solution of the BEQs (2.9) and (2.12). During (re)heating, the inflaton energy density scales as $\rho_{\phi}(a)\propto a^{-4}$, while the radiation energy density scales as $\rho_{R}(a)\propto a^{-3}$ in the case of fermionic decays, or as $\rho_{R}(a)\propto a^{-1}$ for scalar decays.
Finally, we note that beyond matter production, the (re)heating era is also a source of a potentially observable stochastic gravitational wave (GW) background, from channels that include: (i) graviton Bremsstrahlung [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94] and graviton pair production from inflaton annihilations [95, 96, 97, 98], (ii) graviton emission from inflaton scattering with either thermalized [99] or non-thermalized [100] decay products, and (iii) scattering among thermalized particles during reheating [101]. For GW production in the radiation-dominated era, see also [102, 103, 104, 105, 106, 107, 108, 109] for studies involving thermal plasma interactions.
3 Here be dragons
Figure 3: Sketch of the typical stages of post-inflationary (re)heating in oscillatory models of inflation, and the approximate computational regimes.
Almost three decades ago, Kofman, Linde, and Starobinsky [110, 111] (see also Refs. [112, 113, 64, 114]) identified fundamental shortcomings in the widely used Boltzmann approach presented in the previous section, demonstrating that it is not just inaccurate, but also fundamentally inadequate, at least during the initial stages of (re)heating; cf. Fig. 3. First, Eq. (2.2) violates the fluctuation-dissipation theorem, which states that dissipation within a system necessarily generates fluctuations [115, 116]. In particular, the analysis does not account for how these fluctuations modify the effective mass of the inflaton condensate. Another issue is that the perturbative approximation employed when computing amplitudes and decay rates neglects both the coherent nature of the inflaton field and crucial phenomena such as Bose enhancement and Pauli blocking effects. Most critically, even for coupling values where radiative corrections still remain negligible, perturbative methods break down. For example, if the inflaton oscillates in a potential steeper than quadratic, homogeneous oscillations can trigger explosive instabilities leading to rapid amplification of field fluctuations. Unlike the slow, steady energy leakage of perturbative decays, this mechanism violently disrupts the inflaton condensate, causing its fragmentation and initiating a highly non-linear interplay with its decay products, which can also undergo non-perturbative enhancements, further complicating the dynamics and often requiring the use of classical lattice simulations for a proper treatment; see e.g. Refs. [117, 118, 119]. [^3] For a justification of the validity of the classical analysis and its limitations, see e.g. Refs. [120, 121, 122]. This section explores these dramatic, non-perturbative phenomena that challenge the conventional perturbative picture presented in Section 2.
3.1 Inflaton self-resonance and fragmentation
At the end of inflation, the inflaton field undergoes coherent oscillations around the minimum of its potential. In models where the potential is of the form (2.1) with $n>2$, these oscillations can trigger self-resonance instabilities, driving the energy transfer from the homogeneous condensate $\bar{\phi}(t)$ into spatial gradients and ultimately leading to its fragmentation. This instability arises because the oscillating field sources its own perturbations $\delta\phi(t,\vec{x})=\phi(t,\vec{x})-\bar{\phi}(t)$ amplifying them non-perturbatively [123, 124, 125, 126, 127, 128]. In Fourier space, we have
${\delta\ddot{\phi}}_{\vec{k}}+3\,H\,\delta\dot{\phi}_{\vec{k}}+\left[\frac{k^{2}}{a^{2}}+V_{,\phi\phi}(\bar{\phi})\right]\delta\phi_{\vec{k}}=0\,,$ (3.1)
with $V_{,\phi\phi}(\bar{\phi})$ a periodic function of time. The general solution to this equation, incorporating the expansion, takes the form [129]
$\delta\phi_{k}=\mathcal{P}_{k+}(t)\,e^{+\mu_{k}\,t}+\mathcal{P}_{k-}(t)\,e^{-\mu_{k}\,t},$ (3.2)
with $\mathcal{P}_{k\pm}(t)$ periodic functions of time determined by the initial conditions, and $\mu_{k}$ the so-called Floquet exponents, which can change with time due to the presence of the expansion term. If the real part of these exponents does not vanish for specific $k$ values and $\operatorname{Re}(\mu_{k})\gg H$, then an unstable solution grows exponentially with time, indicating non-adiabatic particle production. If efficient enough, this growth of perturbations eventually results in a backreaction on the condensate. As a result, the initially homogeneous field breaks into localized, transient structures, accelerating the energy transfer away from the homogeneous mode. At this stage, particle occupation numbers become poorly defined, and a nonlinear wave description is more suitable [123, 124, 125, 126].
Although the (pre)heating dynamics is often thought to be primarily dictated by the shape of the potential near its minimum, surprising effects can emerge when the potential becomes shallower than quadratic at large field values. In such cases, fragmentation can give rise to localized, long-lived structures known as oscillons [130, 131, 132]: pseudostable, non-topological solitons that emerge from the nonlinear dynamics of the inflaton field, cf. Fig. 4. For instance, while the condensate does not fragment in $T$-models of inflation with $n=2$ and $\Lambda\sim M_{P}$, [^4] Note that this implicitly disregards gravitational effects, which have been shown to ultimately lead to the formation of nonlinear structures, even in free field cases, much like the gravitational instabilities seen in pressureless matter in the late Universe [133, 134]. long-lived oscillon configurations form efficiently for $\Lambda\ll M_{P}$, collectively acting as dust and dominating the energy background for many $e$-folds of post-inflationary expansion [135] (see also Refs. [136, 137]). However, for potentials steeper than quadratic ($n>2$), oscillons are generally not formed, and fragmentation leads to transient structures that decay more rapidly [135, 138], giving rise to a radiation-dominated Universe.
Figure 4: Evolution of the inflaton energy overdensities during oscillon formation in Einstein-Cartan Higgs inflation (adapted from Ref. [137]). The snapshots, extracted from 3+1 classical lattice simulations, display regions where the local energy density exceeds the average by factors of 6 (light blue) and 20 (dark blue) at various times after inflation. Early-time fragmentation generates transient overdensities, while true oscillons—localized, quasi-spherical configurations—emerge by $M\,t\simeq 500$, with $M^{-1}$ denoting the characteristic inflaton oscillation timescale.
A key consequence of fragmentation is its impact on the inflaton equation of state describing the ratio of the spatially averaged pressure $\langle p_{\phi}\rangle$ to the energy density $\langle\rho_{\phi}\rangle$. For any virialized late-time configuration [$1/2\langle{\dot{\phi}^{2}}\rangle=1/2\langle{(\nabla\phi/a)^{2}}\rangle+n\langle{V}\rangle$] involving a monomial potential of the form $V(\phi)\sim|\phi|^{n}$, we have
$\langle w_{\phi}\rangle\equiv\frac{\langle p_{\phi}\rangle}{\langle\rho_{\phi}\rangle}=\frac{\langle\dot{\phi}^{2}/2-(\nabla\phi)^{2}/6a^{2}-V\rangle}{\langle\dot{\phi}^{2}/2+(\nabla\phi)^{2}/2a^{2}+V\rangle}=\frac{1}{3}+\frac{2}{3}\frac{n-4}{(n+2)+2\langle(\nabla\phi/a)^{2}\rangle/\langle V\rangle}\,.$ (3.3)
Note that, unlike Eq. (2.11), this expression explicitly incorporates contributions from the gradient energy density, capturing the full nonlinear evolution of the inflaton field. Initially, when the gradients are small $\langle(\nabla\phi/a)^{2}\rangle\ll\langle V\rangle$, the equation of state is well approximated by the homogeneous field result in that expression, for all values of $n$. However, the late evolution is sensitive to the details of the potential [65, 66]. For $n=2$ and $\Lambda\ll M_{P}$, oscillons form, and $w_{\phi}$ remains close to zero. In contrast, for $n>2$, fragmentation leads to strong interactions among inflaton fragments, driving rapid energy redistribution regardless of the value of $\Lambda$. This process results in a radiation-dominated phase of expansion, $\omega_{\phi}\simeq 1/3$, for $\langle(\nabla\phi/a)^{2}\rangle\gg\langle V\rangle$ [139, 135, 140].
3.2 Stimulated boson production
Beyond the self-interactions of the inflaton field discussed in the previous sections, non-perturbative effects can also play a crucial role in the dynamics of additional matter degrees of freedom coupled to the inflaton. In particular, bosonic fields that interact directly with the inflaton can experience efficient particle production through mechanisms such as parametric [112, 113, 64, 110, 111] or tachyonic resonance [141, 142, 143], triggered by coherent oscillations of the inflaton condensate after the end of inflation. To understand the early evolution of quantum fluctuations during this stage, one can consider the linearized equation of motion for a scalar field $\varphi$, coupled to an oscillating background inflaton $\phi(t)$ in an expanding Friedmann–Lemaître–Robertson–Walker Universe. The Fourier modes of the perturbations $\delta\varphi_{\vec{k}}$ obey the general equation
$\delta\ddot{\varphi}_{\vec{k}}+3\,H\,\delta\dot{\varphi}_{\vec{k}}+\left[\frac{k^{2}}{a^{2}}+{\cal M}_{\varphi}^{2}(\bar{\phi})\right]\delta\varphi_{\vec{k}}=0\,,$ (3.4)
with the effective mass term ${\cal M}_{\varphi}^{2}$ depending on the background inflaton field $\bar{\phi}(t)$. This general form serves as a starting point for analyzing a wide variety of (pre)heating scenarios. In particular, in the presence of three-legged vertex scalar interactions $\mu\,\phi\,|\varphi|^{2}$, the daughter particles $\varphi$ acquire an effective mass ${\cal M}_{\varphi}^{2}=\mu\,\phi(t)$ due to the oscillating inflaton condensate. The trilinear vertex therefore results in a tachyonic mass of $\varphi$, whenever $\phi(t)<0$. Consequently, the modes satisfying $k^{2}/a^{2}<\mu\,|\phi|$ will be exponentially amplified during a portion of each half-period of the inflaton oscillations. This trilinear vertex therefore gives rise to both parametric resonance [64, 114, 110, 111] and tachyonic (pre)heating [141, 142], leading to what is known as tachyonic resonance [144, 143, 145]. This results in very efficient production of the daughter particles, which is not captured by the perturbative calculations. If the amplitude $\Phi$ of inflaton oscillations satisfies $\mu\,\Phi>m_{\phi}^{2}$ in the trilinear case, high-order Feynman diagrams give comparable predictions to the lowest-order ones and the problem has to be approached non-perturbatively. Similarly to the case of the three-legged vertex, in the four-legged vertex it is also possible to realize the effect of parametric resonance. In particular, if the effective mass of the daughter field ${\cal M}_{\varphi}^{2}=2\,g\,\phi^{2}(t)$ is much greater than that of the inflaton inflaton, $\sqrt{2\,g}\,\Phi/m_{\phi}\gg 1$, the corresponding frequency in Eq. (3.4) becomes much higher than that of the inflaton, leading to a broad parametric resonance regime whose early stages can be studied using a WKB approximation [110, 111]. In this scenario, the resonant production of the particles occurs over a broad range of $k$, and the (re)heating becomes extremely efficient.
The homogeneous oscillations of the inflaton leading to rapid growth of spatially varying perturbations via parametric or tachyonic resonance cannot proceed forever. As the occupation numbers of the produced particles grow, their collective energy density becomes significant enough to influence the dynamics of the inflaton field [146, 147, 148]. This backreaction effect not only alters the effective potential of the inflaton but also modifies the resonance structure inherent in parametric amplification, leading to the shift or suppression of resonance bands. The simplest way to take into account the backreaction of the amplified quantum fluctuations is to use the Hartree approximation [110, 111], in which different modes and fields evolve independently (uncorrelated in time). However, as the number of particles increases, the mean-field/Hartree approximation stops being a good description, and coupling between different Fourier modes becomes important, necessitating a fully non-linear treatment to accurately capture the evolution of the system. Moreover, the produced particles interact among themselves and with the inflaton field through rescattering processes. These interactions lead to a redistribution of energy among different field modes, contributing to the fragmentation of the inflaton condensate and the approach toward thermal equilibrium.
3.3 Multifield blocking effects
The picture presented in the previous section may change significantly in the presence of multiple scalar fields $\phi_{I}$ that contribute to inflation. In such multifield setups, each inflaton component typically oscillates with a different mass and amplitude, leading to a superposition of oscillatory modes. If all the fields are coupled to a common matter species—for instance, through interaction terms like $\sum_{I}g_{I}^{2}\,\phi_{I}^{2}\,\varphi^{2}$—the resulting effective mass of the daughter field $\varphi$ becomes a time-dependent function with contributions from all $\phi_{I}(t)$ components. Due to the differing frequencies, phases, and decay rates of these inflaton oscillations, the combined evolution often lacks the regularity required for efficient resonance. This loss of coherence between the inflaton fields—sometimes referred to as de-phasing—can “block” the amplification of the daughter field’s fluctuations by smoothing out the zero crossings of its effective mass. As a result, sharp adiabaticity violations necessary for triggering explosive particle production may be suppressed. This phenomenon reduces the efficiency of (pre)heating, delaying energy transfer to the radiation bath, and potentially prolonging the (re)heating phase.
Despite the aforementioned suppression, multifield models are not necessarily insensitive to non-perturbative processes. The structure of the couplings plays a critical role. For example, introducing trilinear couplings or allowing negative couplings for certain inflaton components can restore or even enhance the resonant instabilities, bypassing the blocking effects associated with simple quadratic couplings.
3.4 The depleting role of fermions
Fermions can also be produced non-perturbatively from the inflaton field during (pre)heating, but their production is severely constrained by the Pauli exclusion principle, which limits their number and reduces their direct impact on the inflaton field [149, 150, 151, 152]. However, in models where resonantly produced bosons decay into fermions, these fermions act as a “spillway” draining energy from the system and alleviating backreaction on the inflaton. This delays the onset of parametric resonance effects, providing a more controlled and gradual evolution of the system [153, 154, 155]. Unlike traditional (pre)heating models, where backreaction typically halts further energy release from the inflaton, this combined (pre)heating mechanism facilitates up to a four-order magnitude increase in energy dissipation, with the remaining inflaton energy reduced to as little as $0.01\%$ [156, 157].
The equation of state of the Universe during (pre)heating is also influenced by the presence of fermions. In conventional (pre)heating models, the energy density evolves from a mixture of matter and radiation. However, in combined (pre)heating, the rapid decay of fermions accelerates the transition to a radiation-dominated state, making the evolution of the Universe faster and more efficient. As fermions act as radiation-like components, they contribute to the overall energy budget, leading to a quicker approach to the characteristic value $w=1/3$ of radiation domination [157].
3.5 Gravitational (pre)heating
The non-perturbative transfer of energy from the inflaton to matter fields can also occur via gravitational interactions, provided that a scalar spectator field $\varphi$—either from the SM or beyond—is non-minimally coupled to gravity, such that its initial dynamics is governed by the effective equation of motion (3.4) with ${\cal M}^{2}_{\varphi}=\xi\,R$, where $\xi$ is the non-minimal coupling to gravity. A well-motivated setting for this mechanism arises in quintessential inflation scenarios, where the Universe undergoes a kination phase ($w_{\phi}=1$) following the end of inflation. During this transition, the Ricci scalar, $R=6\,(\dot{H}+2\,H^{2})=3\,(1-3\,w_{\phi})\,H^{2}$, becomes negative and induces an effective tachyonic mass ${\cal M}^{2}_{\varphi}<0$ for the spectator field $\varphi$. This results in a burst of nonadiabatic particle production via tachyonic instability, leading to the exponential amplification of quantum fluctuations [158, 140, 159]. As first pointed out in Refs. [141, 142, 158], any perturbative picture of a homogeneously oscillating scalar field fails to capture the full dynamics of spontaneous symmetry breaking. In particular, this simplified view also overlooks crucial effects such as the formation of topological defects [160, 158, 140], which can have significant observational consequences, including the generation of a stochastic GW background [160, 161] and baryogenesis [162, 163].
As shown in Refs. [164, 165], the explosive particle production during a Hubble-induced tachyonic phase allows the Higgs field itself to be responsible for the onset of the hot Big Bang era, allowing for (re)heating temperatures in the range $10^{-2}-10^{9}$ GeV and opening the gate to implement potential electroweak baryogenesis mechanisms [166, 167]. The viable parameter space, particularly the relationship between the Higgs mass, the inflationary scale, and the top-quark mass, is constrained by both vacuum stability requirements and the need to achieve successful (re)heating before BBN [164, 165]. In general, these constraints favor a lower top-quark pole mass, in agreement with current measurements [168, 169].
3.6 Towards thermalization
After the initial stages of (pre)heating, where non-perturbative processes lead to explosive particle production, the Universe transitions toward thermalization—a phase characterized by the establishment of thermal equilibrium among the produced particles. This process is intricate, involving various stages and mechanisms that collectively drive the system from a highly non-equilibrium state to one of local thermal equilibrium.
Immediately after fragmentation, the Universe is populated by a dense assembly of interacting fields with large occupation numbers, rendering a classical field theory description appropriate. The energy spectra of the particles produced during (pre)heating are usually highly non-thermal, exhibiting specific peaks corresponding to resonant modes. As interactions proceed, processes such as scattering and decay lead to a redistribution of energy and the establishment of a turbulent dynamics [123, 124, 125, 126], where energy is transferred across different scales, a phenomenon known as an energy cascade. This turbulent behavior is analogous to wave turbulence observed in other physical systems and plays a pivotal role in the redistribution of energy among the modes of the fields. The particle spectra during a turbulent stage exhibit generically a momentum dependence $n_{k}\sim k^{-3/2}$ and a self-similar evolution, characterized by specific scaling behaviors in the distribution functions of the fields. For instance, in the case of a $\lambda\,\phi^{4}$ model, one has [125, 126]
$n_{k}(\tau)=\tau^{-q}\,n_{0}(k\,\tau^{-p})\,,$ (3.5)
with $\tau=\eta/\eta_{c}$ a rescaled conformal time, $n_{0}(k)$ the distribution function at a later time $\eta_{c}$ within the self-similar regime and $q\sim 3.5\,p$ and $p\sim 1/5$ [125, 126]. The presence of such spectra suggests that the system is undergoing a universal process of energy transfer from small- to large-momentum scales, regardless of the specific details of the initial conditions. Such a self-similar dynamic, characteristic of turbulent Kolmogorov spectra, is crucial for understanding how the system approaches thermal equilibrium over time.
Before reaching complete thermalization, the system may enter a state known as prethermalization where certain macroscopic quantities, such as the equation of state, approach values close to those of thermal equilibrium, even if this is not yet locally established [170, 171, 172, 139]. Specifically, the global equation of state parameter tends toward $1/3$, signaling a radiation-dominated Universe. However, true local thermal equilibrium requires both kinetic equilibrium (a uniform distribution of energy and momentum among particles) and chemical equilibrium, with number-changing processes establishing the equilibrium abundances of different species. Although prethermalization can occur relatively quickly, full thermalization is a more gradual process, governed by the rates of these microscopic interactions. Capturing this final evolution presents a significant challenge for classical lattice simulations, as they are fundamentally limited by the well-known Rayleigh-Jeans divergence [121]. In particular, in classical field theory, all modes are thermally occupied according to the Rayleigh-Jeans law, leading to an ultraviolet divergence in energy density at finite temperature. In the continuum limit, this results in an unphysical scenario where the temperature approaches zero, while on a lattice, it remains sensitive to the imposed discretization scale. As a result, while lattice simulations effectively describe early nonlinear dynamics—such as self-resonance, fragmentation, and turbulent energy cascades—they fail to accurately capture the true quantum thermalization process. To properly account for the final approach to equilibrium, more standard techniques, such as quantum BEQs or statistical field theory approaches, are often required [125, 126, 69].
4 Rare analytic birds
As discussed in the previous sections and summarized in Fig. 3 and Table 1, the treatment of (re)heating is often far more intricate than typically assumed within the particle physics community, frequently requiring advanced numerical techniques that go far beyond the simplistic Boltzmann approach. This naturally raises the question: must one always resort to daunting tools like lattice simulations, spending countless hours poring over evolving field configurations? While the general answer is regrettably affirmative, there do exist exceptional scenarios where analytic control can still be partially retained, allowing for meaningful progress without the full machinery of non-perturbative methods. In this section, we highlight some of these rara avis.
I. Foundational Assumptions
Topic Typically assumed Commonly happening Conceptual Implications
Fields Treated as particle gases. Behave as classical waves for large occupation numb. Standard particle-based tools become unreliable.
Quantum stat. Often ignored. Fermions treated like bosons. Bose effects enhance growth Pauli exclusion limits it. Quantum statistics shapes particle production.
II. Inflaton and Energy Transfer
Inflaton Smooth and oscillating condensate. Breaks into inhomogeneous lumps/solitons/defects. Full field dynamics needed. Lattice methods required.
Energy Transf. Slow, perturbative decays. Expansion neglected. Fast, explosive instabilities. Redshift affects resonances. Nontrivial energy loss. Lattice methods required.
Backreaction Homogeneity unaffected. Nonlinearities ignored. Breaking of homogeneity.   Rescatterings & turbulence. Fully coupled dynamics. Lattice methods required.
III. Spectrum, Timescales and Thermalization
Spectrum Narrow and peaked at production Broad with distinct bands. Spectral structure controls the dynamics.
Timescale Radiation domination onset takes many oscillations. (Re)heating can complete in just a few oscillations. Transition may be nearly instantaneous.
Thermalization Instant once energy is transferred. Delayed by chaotic field behavior. Needs modeling of prethermal stages.
Final State Equilibrium bath with clear temperature. Chaotic fields far from equilibrium. Must distinguish early and final temperatures.
Table 1: Comparison between the Boltzmann ((re)heating) and non-perturbative ((pre)heating) approaches.
4.1 Starobinsky scenario: a perturbative take
Among the rare instances where a perturbative treatment of (re)heating remains reliable, the Starobinsky model of inflation stands out as a prime example. The standard action of this scenario takes the form [38, 39, 40, 41, 173]
$S=\frac{M_{P}^{2}}{2}\int d^{4}x\,\sqrt{-\tilde{g}}\left[\tilde{R}+\frac{\tilde{R}^{2}}{6M^{2}}\right]+S_{M}(\tilde{g}_{\mu\nu},\varphi)\,,$ (4.1)
with $S_{M}$ a matter action containing all SM model fields, and in particular, the Higgs doublet $\varphi$ minimally coupled to gravity. The presence of the term $\tilde{R}^{2}$ in Eq. (4.1) allows an inflationary state able to generate the observed amount of primordial density perturbations ${\cal P}\simeq 2.1\times 10^{-9}$ [37] for suitable values of the mass parameter $M$, namely $M\simeq 1.3\times 10^{-5}\left(54/N_{*}\right)M_{P}$. Here, $N_{*}$ corresponds to the number of $e$-folds of inflation needed to solve the flatness and horizon problems, which depends itself on the whole post-inflationary history and, in particular, on the details of the heating stage, which we now describe.
Unlike other approaches that require introducing couplings between the inflaton and matter fields by hand, the Starobinsky scenario offers a predictive and universal framework where all interactions arise unambiguously from gravitational couplings. In particular, once the theory is defined in the frame (4.1) with a non-minimally coupled matter sector, no further assumptions are needed: no ad hoc interaction terms, arbitrary branching ratios, or speculative particle content. To see this explicitly, one can perform a Weyl transformation $\tilde{g}_{\mu\nu}=\Omega^{2}(\phi)\,g_{\mu\nu}$ with conformal factor $\Omega^{2}=\exp(\sqrt{2/3}\,\phi/M_{P})$, such that the action is recast as [174, 175, 176]
$S=\int d^{4}x\sqrt{-g}\left[\frac{M_{P}^{2}}{2}R-\frac{1}{2}\,g^{\mu\nu}\,\partial_{\mu}\phi\,\partial_{\nu}\phi-\frac{3}{4}\,M_{P}^{2}\,M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}\right)^{2}\right]+S_{M}(\Omega^{2}(\phi)g_{\mu\nu},\varphi)\,,$ (4.2)
with the so-called scalaron field $\phi$ playing the role of the inflaton field. In this Einstein or scalaron frame, the action for the nonconformally coupled Higgs boson inherits universal couplings to the scalaron through the conformal rescaling of the metric. Crucially, the rescaled field-dependent mass term in this expression never vanishes or becomes negative during the post-inflationary evolution, preventing the emergence of tachyonic instabilities or resonance bands that would otherwise trigger non-perturbative particle production. In addition, the formation of massive oscillons, typically associated with potentials that away from the minimum are shallower than quadratic, does not take place in this setting since $\Lambda\sim M_{P}$ [177].
The absence of non-perturbative phenomena greatly simplifies the post-inflationary dynamics. The scalaron decays perturbatively into the Higgs field via gravitationally induced interactions, whose strength is fixed by the model and not subject to tuning. [^5] We note that, although absent at tree level, the anomalous decay of the scalaron field into gauge bosons is possible at 1-loop. In particular, the breaking of scale symmetry during the regularization process translates into an induced breaking of the gauge conformal symmetry [178]. The resulting decay width can be computed analytically via the Bogoliubov’s method in the original frame (4.1[38, 39] or through standard perturbative techniques in the scalaron frame (4.2[179, 178], obtaining in both cases [180, 96]
$\Gamma_{\phi}\simeq\frac{1}{24\pi}\frac{M^{3}}{M_{P}^{2}}+\mathcal{O}\left(\frac{m_{h}}{M}\right)^{2}\simeq 2.9\times 10^{-17}\,M_{P}\,,$ (4.3)
where we have taken into account that the Higgs mass $m_{h}$ is much smaller than the scalaron mass, neglecting therefore phase-space suppression factors. Similarly, assuming instantaneous thermalization of the decay products, [^6] Strictly speaking, the decay products are initially distributed with smaller occupation numbers and harder momenta, $\langle k\rangle\simeq M$, compared to those in a thermal distribution [70, 181, 182]. As a result, SM particles become approximately thermalized only at a later stage, characterized by the temperature $T_{\rm th}\simeq\alpha^{4/5}M\left(\Gamma_{\phi}\,M_{P}^{2}/M^{3}\right)^{2/5}$, with $\alpha$ being the fine structure constant of the gauge interaction [70]. For a minimally coupled Higgs and $\alpha\simeq 10^{-2}$, this corresponds to $T_{\rm th}\simeq 10^{11}$ GeV; about one order of magnitude below $T_{\text{max}}$. Since $T_{\text{max}}$ depends more weakly on $\Gamma_{\phi}$ than does $T_{\rm th}$, the thermalization temperature can exceed the maximum temperature, $T_{\rm th}>T_{\text{max}}$, if $\Gamma_{\phi}>2\times 10^{-9}M_{P}$, which justifies the assumption of instantaneous thermalization. On the other hand, if $T_{\text{max}}\gtrsim T_{\rm th}$, then $T_{\text{max}}$ should be interpreted as an effective measure of the maximal radiation density as defined by Eq. (4.4). the maximal effective radiation temperature of the associated SM plasma can be estimated as
$T_{\text{max}}=\left(\frac{30\,\rho_{R}(T_{\text{max}})}{\pi^{2}\,g_{\star}}\right)^{1/4}\simeq\left(\frac{5\sqrt{3}}{\pi^{2}\,g_{\star}}\sqrt{\,V}\,M_{P}\,\Gamma_{\phi}\right)^{1/4}\simeq 2\times 10^{12}~{}\text{GeV}\,.$ (4.4)
The temperature of the SM plasma within this period is therefore in the range $T_{\text{rh}}\lesssim T\lesssim T_{\text{max}}$, with $T_{\text{rh}}\simeq g_{\star}^{-1/4}\sqrt{\Gamma_{\phi}\,M_{P}}\simeq 4.2\times 10^{9}$ GeV the standard (re)heating temperature, defined approximately by the moment at which the total decay width of the scalaron into the SM components equals the Hubble rate, $\Gamma_{\phi}=3\,H$ [183, 184].
4.2 Non-oscillatory potentials: an analytical non-perturbative approach
In contrast to standard (pre)heating scenarios that require prolonged oscillations of the inflaton and lattice simulations to follow nonlinear evolution, a particularly elegant and analytically tractable mechanism known as instant (pre)heating emerges in non-oscillatory models [185, 186]. In these scenarios, the inflaton field does not oscillate after inflation but rolls monotonically, enabling non-perturbative particle production through a single, sharp violation of adiabaticity. This one-time event allows the entire mechanism to be described analytically, bypassing the need for nonlinear numerical tools. For this to happen, the effective mass term ${\cal M}_{\varphi}^{2}$ for the matter field $\varphi$ should be $i)$ large enough during inflation as to retain the single-field inflationary dynamics, $ii)$ vary rapidly at the end of inflation to heat the Universe via adiabaticity violations, and, depending on the model, $iii)$ decrease monotonically with time in order to avoid strong backreaction effects at large $\phi$ values. A simple choice satisfying all these criteria is, for instance, [^7] Alternative options that share the characteristics described in $i)$, $ii)$, and $iii)$ could be used without significantly modifying the conclusions below. For example, one could introduce a parameter $\phi_{l}$ encoding the timing of the transition simply by replacing $\phi$ by $\phi-\phi_{l}$ or one could consider smoothing the transition at $\phi=0$ by using some interpolation function; see, e.g. Ref. [187]. Note also that condition $iii)$ could be relaxed if the daughter field $\varphi$ is coupled to lighter fermionic degrees of freedom. Their inclusion significantly increases the efficiency of (re)heating, draining energy from the $\varphi$ sector before the backreaction on $\phi$ becomes important [185, 186]. In the most efficient scenarios, (re)heating can become nearly instantaneous.
${\cal M}_{\varphi}^{2}(\phi)=\begin{dcases}g^{2}\,\phi^{2}&\quad\text{ for }\phi\leq 0\,,\\ \tilde{m}^{2}_{\varphi}&\quad\text{ for }\phi>0\,,\end{dcases}$ (4.5)
with $\tilde{m}_{\varphi}$ a constant. This a priori unconventional behavior is expected in models where quintessential inflation is associated with the emergence of quantum scale symmetry in the vicinity of UV and IR fixed points [188, 187, 189]. In this type of variable gravity scenario [190], rapid variations of ${\cal M}_{\varphi}^{2}(\phi)$ are only expected to occur in a crossover regime where the dimensionless couplings and mass ratios of matter fields evolve from their UV to IR values. If the field $\varphi$ is identified with the SM Higgs, the decoupling from $\phi$ in Eq. (4.5) at late times encodes the approach to the SM IR fixed point, as required by the constraints on the variation of the Fermi to Planck mass ratio since nucleosynthesis [191, 192].
The form of Eq. (4.5) ensures that ${\cal M}_{\varphi}^{2}$ changes non-adiabatically only once near $\phi=0$, triggering an instantaneous burst of particle production. At small $k$ values, the violation of the adiabaticity condition can be safely approximated by |˙$|\dot{\cal M}_{\varphi}|\gtrsim{\cal M}^{2}_{\varphi}$ or equivalently by $g^{2}\,\phi^{2}\lesssim g\,|\dot{\phi}_{0}|$, with $|\dot{\phi}_{0}|$ the inflaton velocity at zero crossing. Solving this expression for $\phi$, we observe that particle production takes place in a very narrow interval $\Delta\phi\sim(|\dot{\phi}_{0}|/g)^{1/2}$ around $\phi=0$, the production being essentially instantaneous for sufficiently large couplings, $\Delta t\sim\Phi/|\dot{\phi}_{0}|\sim(g\,|\dot{\phi}_{0}|)^{-1/2}$. The typical momentum of the created particles follows directly from the uncertainty principle, $\Delta k\sim(\Delta t)^{-1}\sim(g\,|\dot{\phi}_{0}|)^{1/2}$ and coincides with the one obtained by properly solving the mode equation (3.4) in the WKB approximation. Indeed, as shown explicitly in the seminal paper [111], the occupation number of $\varphi$ particles after a single zero crossing is given by
$n^{\rm kin}_{k}=\exp\left(-\frac{\pi\,k^{2}}{g\,|\dot{\phi}_{0}|}\right).$ (4.6)
Assuming the decay products to be ultra-relativistic ($g\,|\dot{\phi}_{0}|\gg\tilde{m}_{\varphi}$), this corresponds to an instantaneous generation of the radiation energy density $\rho^{\rm kin}_{R}\simeq g^{2}/(4\pi^{4})\,|\dot{\phi}_{0}|^{2}$, which redshifts as $\rho_{R}\propto a^{-4}$. Since no further non-adiabatic transitions occur, there is no subsequent resonance, and the entire process is completed in a single event, in stark contrast with oscillatory models where repeated zero-crossings lead to parametric resonance and require careful numerical treatment. On top of that, the results are completely independent of the particle spin, allowing us to extend the above estimates to fermionic species. We emphasize that this is not the case in oscillatory scenarios, since the adiabaticity condition is violated periodically, leading to bosonic enhancement effects [110, 111].
Using the scaling of the different energy components during the kination epoch ($\rho_{\phi}(a)\propto a^{-6}$ and $\rho_{R}(a)\propto a^{-4}$) together with entropy conservation, one can easily determine the (re)heating temperature at which the energy density of the created particles equals that of the inflaton field ($\rho_{R}(T_{\text{rh}})=\rho_{\phi}(T_{\text{rh}})$), namely
$T_{\text{rh}}=\left(\frac{g_{\star s}(T_{\text{max}})}{g_{\star s}(T_{\text{rh}})}\right)^{1/3}\Theta^{1/2}\,T_{\text{max}}\,,$ (4.7)
with $g_{\star s}(T_{\text{max}})$ and $g_{\star s}(T_{\text{rh}})$ the entropic degrees of freedom at the corresponding temperature scales, $T_{\text{max}}^{4}\equiv 30\,\rho_{\rm R}^{\rm kin}/({\pi^{2}\,g_{\star}(T_{\text{max}})})$ the maximum temperature of the created particles at the onset of kination under the assumption of instantaneous thermalization, and
$\Theta\equiv\frac{\rho_{R}^{\rm kin}}{\rho_{\phi}^{\rm kin}}\simeq 2\times 10^{-8}\left(\frac{g}{0.02}\right)^{2}\left(\frac{10^{11}\,{\rm GeV}}{H_{\rm kin}}\right)^{2}\left(\frac{|\dot{\phi}_{0}|}{10^{-8}M_{P}^{2}}\right)^{2}$ (4.8)
the so-called heating efficiency encoding the efficiency of non-perturbative particle production [187, 193]. We observe then that the smaller the heating efficiency, the longer the kination epoch, and the larger the difference between the maximal radiation temperature and the proper (re)heating temperature. In particular, accounting for all SM degrees of freedom at temperatures higher than the top-quark mass ($g_{\star}(T_{\text{max}})=g_{\star s}(T_{\text{max}})=g_{\star s}(T_{\text{rh}})=106.75$), the upper limit on the inflationary scale $H_{I}<2.5\times 10^{-5}\,M_{P}$ [37] implies a bound $T_{\text{max}}\leq 6\times 10^{11}~{}\text{GeV}\times\Theta^{1/4}$.
5 Conclusions
As particle physics and cosmology continue to further intertwine, the (re)heating epoch stands out as a pivotal bridge between high-energy inflationary dynamics and the thermal history of the observable Universe. A key takeaway from this review is that, despite its intuitive appeal and analytical tractability, the conventional Boltzmann approach—commonly used within the particle physics community—fails to capture the essential nonlinear and non-perturbative phenomena that dominate the early stages of (re)heating. These include explosive particle production via parametric and tachyonic resonance, condensate fragmentation, and turbulent energy cascades; processes that are not just theoretical curiosities but critical components of the post-inflationary dynamics.
Although significant uncertainties concerning the detailed evolution from the end of inflation to radiation domination remain, decades of theoretical progress and increasingly sophisticated computational tools, ranging from lattice simulations to semi-analytic techniques, have brought us closer to a coherent and predictive framework. One of the aims of this review has been to demystify these developments, highlight the physical processes at play, and offer to the particle physics community a practical entry point into this rich area of early-Universe dynamics. Moreover, the dynamics of (pre)heating is not merely of academic interest: they can leave observable imprints. The violent amplification of field fluctuations can source stochastic gravitational waves [194], contribute to baryogenesis [195, 196, 197, 198, 199, 200, 201], affect the primordial power spectrum [202, 203, 204, 205, 206, 207, 208], and lead to the formation of topological defects such as cosmic strings or domain walls [209, 210, 141, 211, 212, 160, 158, 140]. These cosmological relics offer a unique window into the Universe’s earliest moments and provide strong motivation for connecting early-Universe dynamics with upcoming observational probes.
In summary, (re)heating is not a footnote to inflation, nor a simple perturbative afterthought. It is a rich and intricate phase that demands both conceptual clarity and methodological care. As cosmology becomes an increasingly precise science grounded in data, faithful modeling of this transitional epoch will be essential not only for consistency but also for unlocking new insights into high-energy physics and the fundamental structure of our Universe.
Acknowledgments
BB and NB, known to have previously relied a bit too faithfully on the Boltzmann approach, hereby publicly commit to respecting its limitations; at least until the next draft. NB received funding from the grant PID2023-151418NB-I00 funded by MCIU/AEI/10.13039/ 501100011033/FEDER, UE. JR is supported by a Ramón y Cajal contract of the Spanish Ministry of Science and Innovation with Ref. RYC2020-028870-I. This research was further supported by the project PID2022-139841NB-I00 of MICIU/AEI/10.13039/501100011033 and FEDER, UE. JR thanks Matteo Piani for valuable assistance with oscillon snapshots.